TABLE 2 Sources of Growth for Groups of Countries

Net  Investment Government  Black  Revolutions
Group  Period Convergence  Share  Share  Market  Per  Fitted  Actual
Effect  in Output in Output Premium  Year  Growth  Growth
14 Sub-Saharan 1965-75  -0.2  -0.6  -1.1  -0.2  -0.3  -2.3  -2.8
Africans  1975-85  0.6  -0.6  -1.0  -0.9  -0.3  -2.2  -3.9
slow growers  1985-95  1.7  -0.7  -1.1  -0.7  -0.3  -1.0
5 Latin-  1965-75  -0.9  0.0  0.2  -0.3  -0.1  -1.1  -1.9
American  1975-85  -0.4  -0.3  -0.8  -1.0  -0.1  -2.5  -4.1
slow growers  1985-95  0.2  -0.3  -1.1  -1.6  -0.1  -2.9
4 Sub-Saharan 1965-75  2.2  -0.4  -0.6  0.0  0.2  1.4  3.2
Africans  1975-85  1.0  0.0  -0.8  0.1  0.2  0.4  1.6
fast growers  1985-95  0.4  0.1  -0.8  0.3  0.2  0.1
3 Latin-  1965-75  0.4  0.2  0.2  0.0  0.0  1.0  2.6
American  1975-85  -0.5  0.2  0.2  0.1  0.0  0.0  1.2
fast growers  1985-95  -0.7  0.2  0.3  0.0  0.0  -0.2
8 East-Asian  1965-75  1.6  0.2  0.6  0.3  0.1  2.8  3.1
fast growers  1975-85  0.8  0.6  1.0  0.6  0.1  3.0  4.2
1985-95  -0.1  0.6  1.0  0.7  0.1  2.3
6 OECD  1965-75  0.5  0.8  0.1  0.4  0.1  2.0  1.6
fast growers  1975-85  0.5  0.5  0.1  0.5  0.1  1.7  2.5
1985-95  -0.5  0.5  0.1  0.7  0.1  0.8

NOTE: Each entry shows the average contribution of the indicated variable to the fitted growth rate of real per capita GDP (expressed relative to the sample mean). The contributions are averages for the designed group of countries and time period. The net convergence effect is the combined impact from the initial values of log(GDP), male and female secondary school attainment, and log(life expectancy). The fitted growth rate is the sum of the contributions shown separately. the actaul growth rate refers to the average deviation from the sample mean for the indicated grup of countries and time period. Source: Barro and Lee (1994), Table 6.